HUANG invested $9000 in mutual funds and bonds. the mutual fund earned 11% simnple interest. the bond earned 6% simple interest. at the end of one year, he had earned $815 in interest. how much did he invested each rate?
Answer by mathstutor458(57)

HUANG invested $9000 in mutual funds and bonds.
let HUANG invested in mutual fund is x, so bond investment=9000-x
the mutual fund earned 11% simnple interest=11% of x=11x/100
the bond earned 6% simple interest=6% of x=6(9000-x)/100
at the end of one year, he had earned $815 in interest.
Therefore,11x/100 + 6(9000-x)/100 =815
(11x+6(9000-x))/100 = 815
11x+54000-6x=815 * 100
5x=81500-54000
5x=27500
x=27500/5=5500
HUANG invested in mutual fund is $5500
HUANG invested in bond is 9000-5500= $4500